Determine if the following are in proportion.
 4, 6, 8, 12 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

4, 6, 8, 12
4 / 6 = 2/ 3
8 / 12 = 2 / 3
Hence 4: 6 = 8: 12
∴These are in a proportion

Thanks